# Hodge structures on conformal blocks

### v1

@@ Abstract.
We prove existence and uniqueness of complex Hodge structures on modular functors, under the assumptions that their conformal blocks are semisimple as representations of mapping class groups and that the associated categories are ribbon. These $2$ assumptions are satisfied by modular functors coming from Lie algebras, and it is an open conjecture that they are always satisfied. The proof is based on the non-abelian Hodge correspondence and Ocneanu rigidity. Given a modular functor, we explain how its Hodge numbers fit into a Frobenius algebra and the Chern characters of its Hodge decompositions into a new cohomological field theory (CohFT). In the case of $\mathrm{SU}(2)$ modular functors of level $2$ times an odd number, we give explicit formulas for all Hodge numbers, in any genus $g$.
@@ 1. Introduction
This article is the second in a series of two which aims at constructing and studying Hodge structures on the flat vector bundles (spaces of conformal blocks) associated to modular functors. In this article, which can be read independently from the first one, we prove existence and uniqueness of these Hodge structures under mild hypotheses, and compute Hodge numbers of $\mathrm{SU}(2)$ modular functors.
@@ 1.1. Modular functors
For any modular functor there exists a pair of integers $r,s\geq 1$ such that the $s$-th power of the central element and all $r$-th powers of Dehn twists act trivially in every $\rho_{g}(\underline{\lambda})$. The quotient of $\mathrm{PM\widetilde{o}d}(S_{g}^{n})$ by these elements is the fundamental group of a $\mu_{s}$-gerbe ¯$\overline{\mathcal{M}}_{g}^{n}(r,s)$ over the compact moduli space ¯$\overline{\mathcal{M}}_{g}^{n}(r)$ of genus $g$ $r$-twisted nodal curves with $n$ marked orbifold points of order $r$ (see Remarks 2.22 and 2.36). The coarse space of ¯$\overline{\mathcal{M}}_{g}^{n}(r,s)$ is just the Knudsen-Deligne-Mumford moduli space of nodal curves $\overline{M}_{g,n}$. By the correspondence between representations of the fundamental group and bundles with flat connections, we can alternatively see each $\rho_{g}(\underline{\lambda})$ as a bundle with flat connection called conformal block:
@@ 1.2. Hodge structures
For $g\geq 3$, it is expected that the mapping class group $\mathrm{PMod}(S_{g})$ satisfies Kazhdan’s property $(T)$. In the case where the local system $\mathcal{V}_{g}$ has a unitary Galois conjugate, this would imply its rigidity and thus, by Simpson’s result [23, 4.5] reproduced below, existence of a CVHS on $\mathcal{V}_{g}$. Note that if $\mathcal{V}$ comes from a simple Lie algebra, $\mathcal{V}_{g}$ has a unitary Galois conjugate.
@@
A potential application of Hodge structures on $\mathcal{V}_{g}(\lambda_{1},\dotsc,\lambda_{n})$ is detecting irreducible. Indeed, any gap in the Hodge decomposition yields a direct sum decomposition. Such gaps could be found by computing Hodge numbers. However, in Section 4.3, we prove that for all $\mathrm{SO}(3)$ and $\mathrm{SU}(2)$ modular functors, the Hodge structures on the $\mathcal{V}_{g}(\lambda_{1},\dotsc,\lambda_{n})$ have no gaps. It would be interesting to study gaps in Hodge numbers more generally for modular functors coming from simple Lie algebras (see Section 4.2 below).
@@ 1.3. Existence and uniqueness
Here, by a shift of $E=\bigoplus_{p+q=m}E^{p,q}$, we mean another decomposition $E=\bigoplus_{p+q=m}E^{\prime p,q}$, where $E^{\prime p,q}=E^{p+s,q+t}$ for a pair $(s,t)\in\mathbb{Z}^{2}$ independent of $(p,q)$.
Our main result is an analog of Simpson’s in the case of modular functors. Although the flat bundles $\mathcal{V}_{g}(\underline{\lambda})$ of a modular functor $\mathcal{V}$ may not in general individually be rigid, they do not admit compatible families of deformations (Ocneanu rigidity). Using this fact and non-abelian Hodge theory, we are able to prove that, under $2$ assumptions, Hodge structures on modular functors exist and are unique up to shifts.
@@
Let $\mathcal{V}$ be a complex modular, genus 00 modular or rooted functor, with semisimple quantum representations and such that the associated weakly ribbon category $\mathrm{C}_{\mathcal{V}}$ is ribbon. Then $\mathcal{V}$ admits a complex Hodge structure. Moreover, $\mathrm{Shifts}(\Lambda)=\{\delta:\Lambda\rightarrow\mathbb{Q}^{2}\mid\delta(0)=0\}$ acts transitively on the set of complex Hodge structures on $\mathcal{V}$, where $\Lambda$ denotes the set of colors of $\mathcal{V}$.
@@
The shift $\delta$ acts on the CVHS on $\mathcal{V}_{g}(\underline{\lambda})$ by shifting its bidegrees by $\sum_{i}\delta(\lambda_{i})$. Let us comment on the $2$ assumptions. That any weakly ribbon [^1] In our definition, a weakly ribbon category is semisimple and finite. category is ribbon is to the author’s knowledge an open conjecture (see for example [5, footnote 7 p.5]). We use the assumption that $\mathrm{C}_{\mathcal{V}}$ is ribbon to be able to apply Ocneanu rigidity.
That $\mathcal{V}$ has semisimple quantum representations means that every flat bundle $\mathcal{V}_{g}(\underline{\lambda})$ is semisimple. We assume this to be able to deduce the existence of CVHS from the $\mathbb{C}^{*}$-fixed point property in the non-abelian Hodge correspondence. We conjecture that this assumption is superfluous:
Conjecture 1.1.
Let $\mathcal{V}$ be a modular functor such that the associated weakly ribbon category $\mathrm{C}_{\mathcal{V}}$ is ribbon. Then $\mathcal{V}$ has semisimple quantum representations.
@@
Let $\mathcal{V}$ be a complex modular functor with semisimple quantum representations such that $\mathrm{C}_{\mathcal{V}}$ is ribbon. Then any complex Hodge structure on the rooted functor 222which is a genus 00 structure! associated to $\mathcal{V}$ extends uniquely to the whole modular functor. In particular, knowing genus 00 Hodge numbers is knowing all Hodge numbers.
@@
The idea that Hodge numbers of rooted functors could be organized in an algebra appeared independently in the recent paper [7] of Belkale, Fakhruddin and Mukhopadhyay. See Sections 1.6 and 4.2 for more comments on their paper.
Moreover, given a Hodge structure on a modular functor $\mathcal{V}$, the characteristic classes of the CVHS on the $\mathcal{V}_{g}(\underline{\lambda})$ can be packaged into a cohomological field theory (CohFT) $\omega$, such that the Frobenius algebra associated to the topological part of $\omega$ is $V$. Hence to any modular functor $\mathcal{V}$ we associate a CohFT over the ring $\mathbb{Q}[u,v]_{\mathrm{rat}}$, unique up to the action of $\mathrm{Shifts}(\Lambda)$. If $\mathcal{V}$ is hermitian, under some assumptions on the hermitian form and the Hodge structure, this CohFT is related to the one constructed by Deroin and Marché in [11, 4.3], see Proposition 3.36. We do not expect this CohFT to be related to the one constructed differently from $\mathcal{V}$ by Andersen, Borot and Orantin in [1].
@@ 1.4. Outline of the proof
We sketch the proof of Theorem 3.21 (existence and uniqueness of CVHS) in the case of modular functors. The proofs for genus 00 modular functors and rooted functors are similar.
@@
Now, as above, given a modular functor with semisimple conformal blocks $(\mathcal{V},\nabla)$ to which the correspondence associates the modular functor of Higgs bundle $(\mathcal{V},\theta)$, for each real number $t\in\mathbb{R}_{+}^{*}\setminus\{1\}$ and weight function $w$ (see Definition 3.13), we have maps:
{${{\{\text{isomorphisms }f:(\mathcal{V},\theta)\rightarrow(\mathcal{V},t\theta)\}}}${structures of CVHS on (${{\{\text{structures of CVHS on }(E,\nabla)\text{ with weight function }w\}}}$$\scriptstyle{r}$$\scriptstyle{i}$
@@
To prove existence of a CVHS structure on $(E,\nabla)$, we apply a rigidity result on modular functors that will imply that $(\mathcal{V},\theta)$ and $(\mathcal{V},t\theta)$ must be isomorphic. This result is Ocneanu rigidity for ribbon categories, that we explain now.
@@
To a modular functor $\mathcal{V}$, one can associate a weakly ribbon category $\mathrm{C}_{\mathcal{V}}$ in a functorial and continuous way. In Section 5, we prove that this association is a fully faithful functor $\mathfrak{mod}\rightarrow\mathfrak{wRib}$, using results of Bakalov-Kirillov and Andersen-Ueno.
@@
Ocneanu rigidity implies that a ribbon category has no nontrivial continuous deformation (Corollary 7.5). Let $(\mathcal{V},\nabla)$ be a modular functor with semisimple conformal blocks such that $\mathrm{C}_{\mathcal{V}}$ is ribbon. Using the continuity of both the map $\mathcal{V}\mapsto\mathrm{C}_{\mathcal{V}}$ and the Simpson correspondence, we have that for any $t\in\mathbb{R}_{+}^{*}\setminus\{1\}$, the category $\mathrm{C}_{\mathcal{V}_{t}}$ associated to $(\mathcal{V},t\theta)$ is a continuous deformation of $\mathrm{C}_{\mathcal{V}}$. Hence $\mathrm{C}_{\mathcal{V}_{t}}$ is isomorphic to $\mathrm{C}_{\mathcal{V}}$. By full faithfulness of $\mathfrak{mod}\rightarrow\mathfrak{wRib}$ and the Simpson correspondence, $(\mathcal{V},\theta)$ and $(\mathcal{V},t\theta)$ are isomorphic. Then any choice of isomorphism $f:(\mathcal{V},\theta)\rightarrow(\mathcal{V},t\theta)$ provides a CVHS structure $r(f)$ on $(\mathcal{V},\nabla)$. This proves existence.
@@
Uniqueness is a consequence of Ocneanu rigidity for endofunctors, which says that any continuous deformation of the identity functor of a ribbon category is given by a very simple formula from some map $\delta:\Lambda\rightarrow\mathbb{C}^{*}$ (see Theorem 7.9). Here $\Lambda$ is the set of simple objects in the category, and equals the set of colors of $\mathcal{V}$ when the category is $\mathrm{C}_{\mathcal{V}}$.
@@
Let $\mathcal{V}^{1}$ and $\mathcal{V}^{2}$ be $2$ Hodge structures on $\mathcal{V}$. Using the map $i$ above, for each $t\in\mathbb{R}_{+}^{*}$ the Hodge structures induce isomorphisms $f^{1}_{t}$ and $f^{2}_{t}$ from $(\mathcal{V},\theta)$ to $(\mathcal{V},t\theta)$. Then the maps $(f^{2}_{t})^{-1}\circ(f^{1}_{t})$ form a continuous family of automorphisms of $(\mathcal{V},\theta)$, which induce a continuous family of automorphisms $k_{t}$ of $(\mathcal{V},\nabla)$ and also of the ribbon category $\mathrm{C}_{\mathcal{V}}$. From Ocneanu rigidity for endofunctors we then deduce that there exists $\delta_{t}:\Lambda\rightarrow\mathbb{C}^{*}$ such that $k_{t}$ is on $\mathcal{V}_{g}(\underline{\lambda})$ multiplication by $\prod_{i}\delta_{t}(\lambda_{i})$. Using $r\circ i=\mathrm{id}$ and the definition of $r$ (see Lemma 6.20), we have $\mathcal{V}^{1}=(\pi\circ\log_{t}|\delta_{t}|)\cdot\mathcal{V}^{2}$ where $(\pi\circ\log_{t}|\delta_{t}|)$ is in $\mathrm{Shifts}(\Lambda)$ and $\cdot$" is the action of $\mathrm{Shifts}(\Lambda)$ on the set of complex Hodge structures. This proves uniqueness, i.e. that $\mathrm{Shifts}(\Lambda)$ acts transitively on the set of complex Hodge structures on $\mathcal{V}$.
@@ 1.5. Hodge numbers of $\mathrm{SU}(2)$ modular functors
This description is an extension of that given by Marché in [19, 3] for signatures of $\mathrm{SU}(2)$ modular functors. In their recent paper [7], Belkale, Fakhruddin and Mukhopadhyay independently computed the same rings $V$ by computing coefficients $a_{k}$ and $b_{k}$ in the proof of Theorem 4.7 with a different method (see [7, Th. 1.11]). See Section 1.6 below for more comments on their work.
@@ 1.6. The case of modular functors associated to $\mathfrak{sl}_{n}$
In their work [7], Belkale, Fakhruddin and Mukhopadhyay study Hodge structures on some geometrically constructed rooted functors $\widetilde{\mathcal{V}}_{\mathfrak{g},\zeta}$ associated to a complex finite dimensional simple Lie algebra $\mathfrak{g}$. The conformal blocks of the functor $\widetilde{\mathcal{V}}_{\mathfrak{g},\zeta}$ coincide, where they are defined (i.e. in genus 00), with those of the usual modular functor $\mathcal{V}_{\mathfrak{g},\zeta}$ associated to $\mathfrak{g}$ (see [4, chp. 7] for a definition). Moreover, in [7, 11.1], they explain how to compute the algebra organizing Hodge numbers of $\widetilde{\mathcal{V}}_{\mathfrak{sl}_{n},\zeta}$ for every $n\geq 2$.
If one is able to prove that actually the framed rooted functors of $\mathcal{V}_{\mathfrak{g},\zeta}$ coincides with $\widetilde{\mathcal{V}}_{\mathfrak{g},\zeta}$ (4.10), then using the main result of this paper, the computations of Belkale, Fakhruddin and Mukhopadhyay will extend painlessly to higher genus, without the need for a geometric model (Proposition 4.11). See Section 4.2 for more details.

### v2

@@ Abstract.
We prove existence and uniqueness of complex Hodge structures on modular functors. The proof is based on the non-Abelian Hodge correspondence and Ocneanu rigidity. Given a modular functor, we explain how its Hodge numbers fit into a Frobenius algebra and the Chern characters of its Hodge decompositions into a new cohomological field theory (CohFT). In the case of $\mathrm{SU}(2)$ modular functors of level $2$ times an odd number, we give explicit formulas for all Hodge numbers, in any genus $g$.
@@ 1. Introduction
This article is the second in a series of two which aims at constructing and studying Hodge structures on the flat vector bundles (spaces of conformal blocks) associated to modular functors. In this article, which can be read independently from the first one, we prove existence and uniqueness of these Hodge structures, and compute Hodge numbers of $\mathrm{SU}(2)$ modular functors.
@@ 1.1. Modular functors
For any modular functor there exists a pair of integers $r,s\geq 1$ such that the $s$-th power of the central element and all $r$-th powers of Dehn twists act trivially in every $\rho_{g}(\underline{\lambda})$. The quotient of $\mathrm{PM\widetilde{o}d}(S_{g}^{n})$ by these elements is the fundamental group of a $\mu_{s}$-gerbe ¯$\overline{\mathcal{M}}_{g}^{n}(r,s)$ over the compact moduli space ¯$\overline{\mathcal{M}}_{g}^{n}(r)$ of genus $g$ $r$-twisted nodal curves with $n$ marked orbifold points of order $r$ (see Remarks 2.23 and 2.39). The coarse space of ¯$\overline{\mathcal{M}}_{g}^{n}(r,s)$ is just the Knudsen-Deligne-Mumford moduli space of nodal curves $\overline{M}_{g,n}$. By the correspondence between representations of the fundamental group and bundles with flat connections, we can alternatively see each $\rho_{g}(\underline{\lambda})$ as a bundle with flat connection called conformal block:
@@ 1.2. Hodge structures
For $g\geq 3$, it is expected that the mapping class group $\mathrm{PMod}(S_{g})$ satisfies Kazhdan’s property $(T)$. In the case where the local system $\mathcal{V}_{g}$ has a unitary Galois conjugate, this would imply its rigidity and thus, by Simpson’s result [30, 4.5] reproduced below, existence of a CVHS on $\mathcal{V}_{g}$. Note that if $\mathcal{V}$ comes from a simple Lie algebra, $\mathcal{V}_{g}$ has a unitary Galois conjugate.
@@
A potential application of Hodge structures on $\mathcal{V}_{g}(\lambda_{1},\dotsc,\lambda_{n})$ is detecting irreducibility. Indeed, any gap in the Hodge decomposition yields a direct sum decomposition. Such gaps could be found by computing Hodge numbers. However, in Section 4.3, we prove that for all $\mathrm{SO}(3)$ and $\mathrm{SU}(2)$ modular functors, the Hodge structures on the $\mathcal{V}_{g}(\lambda_{1},\dotsc,\lambda_{n})$ have no gaps. It would be interesting to study gaps in Hodge numbers more generally for modular functors coming from simple Lie algebras (see Section 4.2 below).
@@ 1.3. Existence and uniqueness
Here, by a shift of $E=\bigoplus_{p+q=m}E^{p,q}$, we mean another decomposition $E=\bigoplus_{p+q=m}E^{\prime p,q}$, where $E^{\prime p,q}=E^{p+s,q+t}$ for a pair $(s,t)\in\mathbb{Q}^{2}$ independent of $(p,q)$.
Our main result is an analog of Simpson’s in the case of modular functors. Although the flat bundles $\mathcal{V}_{g}(\underline{\lambda})$ of a modular functor $\mathcal{V}$ may not in general individually be rigid, they do not admit compatible families of deformations (Ocneanu rigidity). Using this fact and non-Abelian Hodge theory we are able to prove that Hodge structures on modular functors exist and are unique up to shifts.
@@
Let $\mathcal{V}$ be a complex modular, genus 00 modular, ribbon or braided functor. Then $\mathcal{V}$ admits a complex Hodge structure. Moreover, $\mathrm{Shifts}(\Lambda)=\{\delta:\Lambda\rightarrow\mathbb{Q}^{2}\mid\delta(0)=0\}$ acts transitively on the set of complex Hodge structures on $\mathcal{V}$, where $\Lambda$ denotes the set of colors of $\mathcal{V}$.
@@
The shift $\delta$ acts on the CVHS on $\mathcal{V}_{g}(\underline{\lambda})$ by shifting its bidegrees by $\sum_{i}\delta(\lambda_{i})$. Note that the notions of modular functors and modular fusion categories are equivalent, thanks to work of Bakalov-Kirillov [4] and a recent paper of Etingof-Penneys [14]. Similarly, the notions of genus 00 modular functor, ribbon functor and ribbon fusion category are equivalent, and those of braided functor and braided fusion category are equivalent. Hence the results above apply to flat bundles associated to modular, ribbon or braided fusion categories.
Remark 1.1.
A flat bundle underlying a CVHS is always semisimple. Hence the statement of Theorem 3.19 contains the semisimplicity of conformal blocks. However, as the proof of semisimplicity relies on results in non-Abelian Hodge theory (formality) which are different from those used in this paper, we have chosen to write it separately in the companion paper [20]. Here are the explicit dependencies between the papers: [20] uses results from Sections 5 and 7; the proof of Theorem 3.19 given in Section 8 relies on [20] and Sections 5, 6 and 7.
@@
Let $\mathcal{V}$ be a complex modular functor. Then any complex Hodge structure on the braided functor 111which is a genus 00 structure! associated to $\mathcal{V}$ extends uniquely to the whole modular functor. In particular, knowing genus 00 Hodge numbers is knowing all Hodge numbers.
@@
The idea that Hodge numbers of braided functors could be organized in an algebra appeared independently in the recent paper [6] of Belkale, Fakhruddin and Mukhopadhyay. See Sections 1.6 and 4.2 for more comments on their paper.
Moreover, given a Hodge structure on a modular functor $\mathcal{V}$, the characteristic classes of the CVHS on the $\mathcal{V}_{g}(\underline{\lambda})$ can be packaged into a cohomological field theory (CohFT) $\omega$, such that the Frobenius algebra associated to the topological part of $\omega$ is $V$. Hence to any modular functor $\mathcal{V}$ we associate a CohFT over the ring $\mathbb{Q}[u,v]_{\mathrm{rat}}$, unique up to the action of $\mathrm{Shifts}(\Lambda)$. If $\mathcal{V}$ is Hermitian, under some assumptions on the Hermitian form and the Hodge structure, this CohFT is related to the one constructed by Deroin and Marché in [11, 4.3], see Proposition 3.34. We do not expect this CohFT to be related to the one constructed differently from $\mathcal{V}$ by Andersen, Borot and Orantin in [1] (see also the related papers by Marian, Oprea, Pandharipande and Zvonkine [26] and by Damiolini, Gibney and Tarasca [10]).
@@ 1.4. Outline of the proof
We sketch the proof of Theorem 3.19 (existence and uniqueness of CVHS) in the case of modular functors. The proofs for genus 00 modular functors, ribbon functors and braided functors are similar.
@@
Now, as above, given a modular functor $(\mathcal{V},\nabla)$ to which the correspondence associates the modular functor of Higgs bundle $(\mathcal{V},\theta)$, for each real number $t\in\mathbb{R}_{+}^{*}\setminus\{1\}$ and weight function $w$ (see Definition 3.11), we have maps:
[…]
@@
To prove existence of a CVHS structure on $(\mathcal{V},\nabla)$, we apply a rigidity result on modular functors that will imply that $(\mathcal{V},\theta)$ and $(\mathcal{V},t\theta)$ must be isomorphic. This result is Ocneanu rigidity for ribbon fusion categories, that we explain now.
@@
To a modular functor $\mathcal{V}$, one can associate a ribbon fusion category $\mathrm{C}_{\mathcal{V}}$ in a functorial and continuous way. In Section 5, we prove that this association is a fully faithful functor $\mathfrak{mod}\rightarrow\mathfrak{Rib}$, using results of Bakalov-Kirillov and Andersen-Ueno.
@@
Ocneanu rigidity implies that a ribbon fusion category has no nontrivial continuous deformation (Corollary 7.5). Let $(\mathcal{V},\nabla)$ be a modular functor with semisimple conformal blocks such that $\mathrm{C}_{\mathcal{V}}$ is ribbon. Using the continuity of both the map $\mathcal{V}\mapsto\mathrm{C}_{\mathcal{V}}$ and the Simpson correspondence, we have that for any $t\in\mathbb{R}_{+}^{*}\setminus\{1\}$, the category $\mathrm{C}_{\mathcal{V}_{t}}$ associated to $(\mathcal{V},t\theta)$ is a continuous deformation of $\mathrm{C}_{\mathcal{V}}$. Hence $\mathrm{C}_{\mathcal{V}_{t}}$ is isomorphic to $\mathrm{C}_{\mathcal{V}}$. By full faithfulness of $\mathfrak{mod}\rightarrow\mathfrak{Rib}$ and the Simpson correspondence, $(\mathcal{V},\theta)$ and $(\mathcal{V},t\theta)$ are isomorphic. Then any choice of isomorphism $f:(\mathcal{V},\theta)\rightarrow(\mathcal{V},t\theta)$ provides a CVHS structure $r(f)$ on $(\mathcal{V},\nabla)$. This proves existence.
@@
Uniqueness is a consequence of Ocneanu rigidity for endofunctors, which says that any continuous deformation of the identity functor of a ribbon fusion category is given by a very simple formula from some map $\delta:\Lambda\rightarrow\mathbb{C}^{*}$ (see Theorem 7.9). Here $\Lambda$ is the set of simple objects in the category, and equals the set of colors of $\mathcal{V}$ when the category is $\mathrm{C}_{\mathcal{V}}$.
@@
Let $\mathcal{V}^{1}$ and $\mathcal{V}^{2}$ be $2$ Hodge structures on $\mathcal{V}$. Using the map $i$ above, for each $t\in\mathbb{R}_{+}^{*}$ the Hodge structures induce isomorphisms $f^{1}_{t}$ and $f^{2}_{t}$ from $(\mathcal{V},\theta)$ to $(\mathcal{V},t\theta)$. Then the maps $(f^{2}_{t})^{-1}\circ(f^{1}_{t})$ form a continuous family of automorphisms of $(\mathcal{V},\theta)$, which induce a continuous family of automorphisms $k_{t}$ of $(\mathcal{V},\nabla)$ and also of the ribbon fusion category $\mathrm{C}_{\mathcal{V}}$. From Ocneanu rigidity for endofunctors we then deduce that there exists $\delta_{t}:\Lambda\rightarrow\mathbb{C}^{*}$ such that $k_{t}$ is on $\mathcal{V}_{g}(\underline{\lambda})$ multiplication by $\prod_{i}\delta_{t}(\lambda_{i})$. Using $r\circ i=\mathrm{id}$ and the definition of $r$ (see Lemma 6.20), we have $\mathcal{V}^{1}=(\pi\circ\log_{t}|\delta_{t}|)\cdot\mathcal{V}^{2}$ where $(\pi\circ\log_{t}|\delta_{t}|)$ is in $\mathrm{Shifts}(\Lambda)$ and $\cdot$" is the action of $\mathrm{Shifts}(\Lambda)$ on the set of complex Hodge structures. This proves uniqueness, i.e. that $\mathrm{Shifts}(\Lambda)$ acts transitively on the set of complex Hodge structures on $\mathcal{V}$.
@@ 1.5. Hodge numbers of $\mathrm{SU}(2)$ modular functors
This description is an extension of that given by Marché in [25, 3] for signatures of $\mathrm{SU}(2)$ modular functors. In their recent paper [6], Belkale, Fakhruddin and Mukhopadhyay independently computed the same rings $V$ by computing coefficients $a_{k}$ and $b_{k}$ in the proof of Theorem 4.7 with a different method (see [6, Th. 1.11]). See Section 1.6 below for more comments on their work.
@@ 1.6. The case of modular functors associated to $\mathfrak{sl}_{n}$
In their work [6], Belkale, Fakhruddin and Mukhopadhyay study Hodge structures on some geometrically constructed braided functors $\widetilde{\mathcal{V}}_{\mathfrak{g},\zeta}$ associated to a complex finite dimensional simple Lie algebra $\mathfrak{g}$. The conformal blocks of the functor $\widetilde{\mathcal{V}}_{\mathfrak{g},\zeta}$ coincide, where they are defined (i.e. in genus 00), with those of the usual modular functor $\mathcal{V}_{\mathfrak{g},\zeta}$ associated to $\mathfrak{g}$ (see [4, chp. 7] for a definition). Moreover, in [6, 11.1], they explain how to compute the algebra organizing Hodge numbers of $\widetilde{\mathcal{V}}_{\mathfrak{sl}_{n},\zeta}$ for every $n\geq 2$.
If one is able to prove that actually the braided functors of $\mathcal{V}_{\mathfrak{g},\zeta}$ coincides with $\widetilde{\mathcal{V}}_{\mathfrak{g},\zeta}$ (4.9), then using the main result of this paper, the computations of Belkale, Fakhruddin and Mukhopadhyay will extend painlessly to higher genus, without the need for a geometric model (Proposition 4.10). See Section 4.2 for more details.